OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4

OpenStudy (aroub):

parallel=same slope y=mx+b , m= -3 , (2,6) 6=-3(2)+b 6=-6+b 12=b so the equation is y=-3x+12